What happens to the force between two charges if the distance between them is (a) halved (b) doubled.

Grade:upto college level

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
The force between two charges is given by
F=\frac{Kq_{1}q_{2}}{r^{2}}
so if the distance is halved
F'=\frac{Kq_{1}q_{2}}{(r/2)^{2}}=4F
so the force increases four times
and, when the distance is doubled
F'=\frac{Kq_{1}q_{2}}{(2r)^{2}}=F/4
and so the force decreases 4 times.